### v2

# SandCell: Sandboxing Rust Beyond Unsafe Code

Abstract.
Rust is a modern systems programming language that ensures memory safety by enforcing ownership and borrowing rules at compile time. While the unsafe keyword allows programmers to bypass these restrictions, it introduces significant risks. Various approaches for isolating unsafe code to protect safe Rust from vulnerabilities have been proposed, yet these methods provide only fixed isolation boundaries and do not accommodate expressive policies that require sandboxing both safe and unsafe code. This paper presents SandCell for flexible and lightweight isolation in Rust by leveraging existing syntactic boundaries. SandCell allows programmers to specify which components to sandbox with minimal annotation effort, enabling fine-grained control over isolation. The system also introduces novel techniques to minimize overhead when transferring data between sandboxes. Our evaluation demonstrates SandCell’s effectiveness in preventing vulnerabilities across various Rust applications while maintaining reasonable performance overheads.
1. Introduction
Rust is a modern system programming language that provides memory-safety by enforcing strict ownership and borrowing rules at compile time. However, in practice, not all Rust programs fully benefit from compiler-enforced memory safety. For example, pre-compiled Rust libraries or libraries written in other programming languages must be called via foreign-function interfaces (FFIs) which the compiler cannot analyze. Also, because the analysis is conservative, some valid, memory-safe Rust programs will fail to type check. Rust provides an unsafe language superset that allows an “escape hatch” from the compiler-enforced safety analysis.
However, unsafe code is generally hard to get right and has become a common source of bugs in Rust (rustfuzztrophy, ). To write correct unsafe code, programmers need to manually maintain the Rust ownership and borrowing rules and other safety invariants expected by the Rust compiler in the unsafe code block. For example, Rust developers must correctly annotate the lifetime of objects beyond the FFI boundary, and ensure any necessary preconditions necessary for memory-safe execution of C functions are met. Failure to meet these expectations can lead to memory-safety vulnerabilities (nistCVE202524898, ; rustsecRUSTSEC20230044Openssl, ).
Even without unsafe code, safe Rust is not immune to memory-safety issues. For example, a long-existing soundness bug (rust-implied-bounds-bug, ) in the type checking of rustc (the official compiler of Rust) allows programmers to arbitrarily expand the lifetime of a reference in safe Rust, which can lead to dangling references. Similarly, the Rust standard library is generally considered part of the safe Rust ecosystem. But given the common use of unsafe code in the standard library, it might contain memory vulnerabilities as well (githubBTreeSetCauses, ; githubFaultPushing, ; githubSoundnessIssue, ).
Given these challenges, there is interest in mitigating lingering memory-safety issues by compartmentalizing Rust programs. A natural idea is to isolate unsafe code from safe code to prevent memory vulnerabilities from affecting the safe subset of the program (almohriFideliusCharmIsolating2018, ; riveraKeepingSafeRust2021, ; kirthPKRUsafeAutomaticallyLocking2022, ; liuSecuringUnsafeRust2020, ; bangTRustCompilationFramework2023, ). These approaches categorize memory objects into safe and unsafe categories according to how they are accessed. Unsafe objects are then placed in a separate compartment, preventing bugs in unsafe code from directly affecting safe objects.
These approaches are, however, limited in several ways. First, their isolation boundary is fixed between safe and unsafe code. In some scenarios discussed in § 3, this boundary is not expressive enough to accommodate fit-for-purpose isolation policies. Second, these approaches do not support separate sandbox instances for different executions of the same piece of code; all executions of the code would execute within one sandbox and share its privileges and runtime data. As a result, they would not support, for example, isolation between handling different requests in a web server, where a separate sandbox is needed for each request, so one request’s handler cannot interfere with the handlers of other requests.
Third, by putting all safe code into the trusted computing base (TCB), all prior approaches are inherently vulnerable to soundness bugs in the rustc as mentioned earlier. Furthermore, prior approaches consider the standard library as part of the TCB, despite the presence of unsafe code in it.
This paper and contributions.
To address these shortcomings of state-of-the-art compartmentalization solutions for Rust, we propose SandCell, a compiler-based approach with three novel properties. Firstly, to allow flexible isolation boundaries while minimizing developers’ effort on annotation, SandCell leverages syntactic boundaries that already exist in the program (e.g., functions and libraries, called “crates” in Rust  [^1] Rust code is distributed as crates, similar to packages or libraries in other languages.) to let developers specify functions, types, or crates as sandboxes. We refer to sandboxes derived from such syntactic boundaries as syntactic sandbox instances. SandCell automatically partitions a Rust application into distinct compartments according to such syntactic instances and inserts sandbox-switching code to enforce the isolation.
To address the second shortcoming, SandCell allows sandboxes to be transient; i.e., every time a certain piece of sandboxed code is invoked, a new run-time instance of the sandbox is created, so that different execution instances of the same code are prevented from accessing each other’s data. This allows SandCell to isolate, for example, the handling of one web request from others requests in a web server.
Finally, SandCell creates isolation boundaries according to developer specifications and does not distinguish between safe and unsafe code. This ensures that the isolation policy defined using SandCell can effectively confine exploits arising from soundness bugs in rustc or memory vulnerabilities in the standard library implementation.
In summary, we make the following contributions:
  • A compiler-based analysis allowing for users to specify syntactic instances of sandboxes for fine-grained and automated sandboxing of Rust functions, types, modules, and entire crates. We discuss the motivation and challenges in § 3 and the design in § 5.
  • Run-time instance sandboxing, which isolates different executions of the same code from each other. We motivate the need in § 3 and describe the design in § 5.
  • SandCell, a general-purpose and highly-automated sandboxing tool for Rust supporting both syntactic instance and run-time instance sandboxing.
  • We evaluate SandCell on several real-world Rust applications, demonstrating its ability to mitigate vulnerabilities and expressing various isolation patterns with reasonable overheads in § 8.
2. Background
This section gives background on Rust, including its memory safety guarantees and memory-unsafe language features; and also discusses the concept of in-process isolation, which is used as our low-level isolation mechanism.
2.1. Memory Safety in Rust
Safe Rust guarantees memory safety by enforcing strict ownership and borrowing rules at compile time. Informally, it has three rules: (1) Every value has a single owner; assignment to a new owner invalidates the previous owner. (2) At any time, a value can be either modified through one reference (either by the owner or through a single mutable reference) or simultaneously read through multiple immutable references, but not both. (3) No reference can outlive the value it points to. To enforce the third rule, Rust associates a lifetime with each reference at the type-system level, which can be viewed as a set of program points where the reference is valid. The Rust compiler can infer the lifetimes of most references, and the rest must be explicitly annotated by the programmer.
The Rust compiler statically enforces these rules to prevent common memory vulnerabilities such as use-after-free and double-free bugs. However, due to the conservative nature of Rust’s static analysis, some valid (i.e., memory-safe) programs might be rejected by Rust’s borrow checker that enforces ownership and borrowing rules at compile time.
2.2. Unsafe Rust
In Rust, the unsafe keyword serves two high-level purposes. First, annotating a function definition with unsafe indicates that the function has contracts that the Rust compiler cannot verify. Second, annotating a function call or code block with unsafe indicates that the programmer has manually ensured the contract is satisfied. In practice, programmers use unsafe code for different reasons, and here we discuss the two most common ones (astrauskasHowProgrammersUse2020, ).
Around 90% of the unsafe code in Rust is used to call unsafe functions (astrauskasHowProgrammersUse2020, ). For example, calling a version of array indexing that does not perform bounds checking is unsafe, since the programmer is responsible for ensuring that the index is within bounds. Another example is calling a function from an external library written in C. In these cases, the developer must ensure that preconditions expected by the unsafe function are satisfied. Violating these preconditions can lead to bugs such as RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ), which is a vulnerability from rust-openssl where a string from Rust code is passed to C code without null-termination, causing an out-of-bound read.
Another common use case of unsafe code is to manipulate raw pointers. It might cause bugs when, for example, the developer assigns an incorrect lifetime to a raw pointer that is cast to a reference. Since the lifetime of a raw pointer is not tracked by the compiler, such bugs cannot be detected at compile time. As a practical example, CVE-2025-24898 (nistCVE202524898, ) is a vulnerability in rust-openssl where an incorrect lifetime parameter is bound to a buffer (allocated in the C code). Because the lifetime parameter exceeds the actual lifetime of the buffer, it leads to a use-after-free vulnerability.
2.3. Bugs in Safe Rust
The official Rust compiler, rustc, is not bug-free either and might accept programs that violate its safety invariants. For example, a well-known bug (rust25860, ) in the type checker enables the lifetime of a local reference to be expanded to the global lifetime. Exploiting such soundness bugs, adversaries may intentionally upload vulnerable crates to crates.io, the Rust package index, to mount supply-chain attacks. One tool Rust developers have at their disposal to minimize the risk posed by third-party dependencies is limiting the dependencies to crates without unsafe code. The package index tags such crates with the “no_unsafe” marker to make finding such crates easier. This mechanism incentivizes malicious developers to exploit soundness bugs instead of unsafe code. The rustc team maintains a list of such compiler bugs (rust-unsound-bug-list, ).
The Rust standard library is also an important building block for safe Rust code as it provides safe abstractions for low-level operations such as heap allocation and synchronization primitives. But bugs in the standard library also exist. For example, an integer overflow led to a buffer overflow in std::str::repeat (RustStdCVE20181000810, ).
2.4. In-process isolation
One way to sandbox cooperating software modules is to place each in a process with its own virtual address space (lamowskiSandcrustAutomaticSandboxing2017, ). However, for tightly-coupled modules, process-based isolation incurs prohibitive context switch and communication overhead (gulmezFriendFoeExploring2023, ). In-process isolation is a security technique that places the code and data for a sandboxed module in its own fault domain, a logically separate portion of the application’s address space. In-process isolation trades strong, hardware- and OS-enforced process isolation for substantially faster context switches between fault domains. In-process isolation can be realized portably and language-independently through software-based fault isolation (SFI) (wahbe1993efficient, ; Tan17, ). SFI, however, increases execution times for the sandboxed modules as their memory accesses must be checked for validity by a software-enforced in-line reference monitor. Modern approaches for in-process isolation leverage hardware features, such as memory protection keys (MPK), to improve upon the efficiency of SFI. The popularity of sandboxed JavaScript in modern browsers, managed languages, and application sandboxing in mobile OSes have motivated processor manufacturers to include such hardware features in commercial, off-the-shelf processors.
Protection Keys for Userspace.
MPK on the x86 architecture uses the protection keys for userspace (PKU) hardware feature. PKU associates every 4KB memory page with a 4-bit protection key identifier which is stored in the OS’s page table entry. A 32-bit CPU register, the protection key rights register (PKRU), controls the access rights of a particular key. The PKRU register holds a 2-bit mask for each of the 16 protection keys. The state of the bits specifies whether pages corresponding to the key are writable/accessible.
In contrast to OS-enforced paging, the PKRU configuration can be modified by the userspace process. Since PKU hardware enforces access control based on the current configuration of the PKRU, unauthorized modification of the PKRU, or techniques that bypass PKU enforcement completely makes PKU vulnerable to threats from run-time attacks and malicious software aiming to escape from its sandbox. For example, an attacker could manipulate /proc/self/mem to bypass MPK protections (connor20, ) or misuse the wrpkru instruction to modify the PKRU register (Vahldiek-Oberwagner19, ).
1
2// Function from C library
3extern C {
4fn parse(req: *const u8) -> *const u8;
5}
[…]
8fn handle_request(
9&self, req: &Socket
10) -> Result<(), Error> {
11let data = req.get_data()?;
12let data = unsafe { parse(req) };
13let key = crypto::get_key()?;
14crypto::verify(&data, &key)?;
15let res = send(&data, req);
16Ok(())
17}
18}
[…]
20let socket = /* omitted */;
21loop {
22let req = accept(socket);
23match handle_request(req) {
24Ok(()) => continue,
25Err(e) => handle_error(e),
26}
27}
28}
[…]
Figure 1. A simplified web server. In Rust, modules (the keyword mod) are used to organize definitions such as functions and types.
3. Motivation and Challenges
Given the disposition of vulnerabilities in unsafe Rust code to undermine the compiler-enforced memory-safety properties, a natural idea is to isolate unsafe code from the rest of a Rust program, as explored by previous work (almohriFideliusCharmIsolating2018, ; riveraKeepingSafeRust2021, ; kirthPKRUsafeAutomaticallyLocking2022, ; liuSecuringUnsafeRust2020, ; bangTRustCompilationFramework2023, ). However, simply having two compartments (one safe and one unsafe compartment) is insufficient for the reasons discussed in § 1 and § 2. Consequently, compartmentalization in Rust must solve the following challenges.
Challenge I: Isolation between unsafe code.
Consider a web server sketched in Figure 1 that handles requests by first parsing the request and then verifying it using a cryptographic library. The parsing and cryptographic libraries are both written in C. In particular, the parse function is directly called via FFIs, while the cryptographic library is wrapped inside a Rust crate crypto.
If we simply put all unsafe code into a separate compartment, a potential memory vulnerability inside the parse function can still be exploited to access the memory of the crypto library, which also contains unsafe code. Therefore, what is more desirable is to put these two libraries into two different sandboxes. This requires the ability to specify isolation boundaries at the library granularity, while existing approaches support only a predefined isolation boundary between all safe and all unsafe code. More generally, the isolation tool should allow developers to isolate different syntactic instances such as functions, modules, and crates.
Challenge II: Isolation between different executions of the same code.
Consider another attack scenario where the adversary sends a malicious request to exploit a use-after-free bug in the crypto library to belonging to other requests that remains in memory. To counter this, a policy isolates the code handling distinct requests from each other is needed. This, however, cannot be accommodated by approaches that sandbox only unsafe code, which shares the states of all requests in a single unsafe-code sandbox. In a broader sense, Challenge II requires the ability to allow the developer to isolate instances of the same code that execute concurrently at run-time from each other.
Challenge III: Isolating soundness bugs and bugs in standard library.
The Rust compiler itself is not bug free and might accept programs that violate its safety invariants, as discussed in § 2.3. By sandboxing untrusted Rust code, the impact of such bugs can be limited to the sandboxed code and data. Such cases requires the ability to isolate a mix of safe and unsafe code.
Rust’s standard library is another important building block for its safety guarantees and provides safe abstractions for low-level operations such as heap allocation and synchronization primitives. Previous approaches that isolate unsafe code from safe code generally include the standard library in the TCB. To mitigate the impact of bugs in the Rust standard library (e.g., (RustStdCVE20181000810, )), the compartmentalization method must support sandboxing not only untrusted and user-specified code, but also dependencies and their transitive dependencies.
Challenge IV: Reducing data exchange cost.
Frequent data exchanges between sandboxes can become a performance bottleneck. The common approach of deep copying data between sandboxes is especially costly as it needs to copy all transitively reachable data.
4. System and Adversary Model
Our system model is similar to other in-process compartmentalization approaches: we assume that the protected application starts in a trusted state from which new, untrusted compartments of sandboxed code can be created at runtime. Similarly to prior work, we also assume that the compartmentalization policy is decided at compile time and remains immutable throughout the lifetime of the application.
We assume an adversary who exploits either pre-existing vulnerabilities in Rust applications, or intentionally introduces vulnerabilities they can exploit into crates that they control. Using such vulnerabilities, the adversary can mount control-flow or data-only attacks against the victim application. The adversary may also attempt to misuse privileges granted to the application, such as performing unwanted, harmful system calls. The goal of SandCell is to mitigate such attacks by isolating untrusted code and data from the rest of the program. However, we do not provide security guarantees against attackers specifically targeting the user-space system-call interpositioning (see  § 7.4). It is also possible that the attacker can indirectly impact the program through information flow between trusted and untrusted code such as through side channels, which is not in the scope of SandCell.
5. Design
In this section, we give an overview of the design of SandCell, including the overall workflow, user specification, its semantics, and the run-time architecture. We will delay the discussion of the program analysis supporting SandCell until § 6 and the implementation details until § 7.
5.1. Overview
SandCell takes the Rust source code and a sandboxing specification as input, and produces a sandboxed binary as output. The SandCell specification contains a list of functions, types, or crates to be sandboxed. We will elaborate on this in § 5.2. SandCell’s analysis builds a call graph of the program and uses the specification to identify the instrumentation points necessary for sandboxing the specified components. SandCell also uses flow-insensitive data-flow analysis to identify allocation sites that might allocate heap objects that are accessed across sandbox boundaries (§ 6). These results are used by the instrumentation component (§ 7.2) of SandCell to alter the program’s source code and insert calls into the low-level isolation application programming interfaces (API) at the instrumentation points identified during the analysis. The in-process isolation APIs and the run-time architecture of SandCell will be described in § 5.4. Finally, the instrumented program is compiled by rustc as normal, producing a sandboxed binary.
5.2. User Specification
1
2[functions] // can be [types] or [crates]
3foo = { transient = true }
4bar = { transient = false }
Figure 2. A SandCell specification example.
In this section, we describe the syntax and semantics of the SandCell specification and how they address the challenges discussed in § 3.
To address Challenge I, SandCell allows users to define sandboxes for functions, modules, crates, and types by listing their names in the specification file, as illustrated by the example in Figure 2, which specifies that functions foo and bar should be sandboxed. When there are multiple candidates with the same name, users can refine it with quantified paths.
In response to Challenge II, we associate a transient field with every sandbox. If transient is set to true, every call to the function will enter a freshly created sandbox. Otherwise, all calls to the function will share the same sandbox instance; i.e., the code and data of the function will be shared across different calls.
For Challenge III, SandCell does not distinguish between safe and unsafe code, but instead follows the user-specified sandboxing boundaries. Therefore, vulnerabilities in the Rust standard library or soundness bugs in rustc can be mitigated by sandboxing the code that uses them.
For Challenge IV, to avoid the high cost of deep copying data between sandboxes, we employ a novel approach that utilizes static analysis to compute allocation sites that may allocate cross-sandbox data, and then instruments those allocation sites so that the data is allocated in some shared data domain between sandboxes that exchange data.
Invocations of different units (functions, crates, etc.) create different sandboxes. For the example in Figure 2, invocations of functions foo and bar create different sandboxes; so their executions are isolated from each other. Conceptually, a sandbox is a memory region that contains the code and data of the sandboxed unit (function, crate, etc.) as well as its transitive dependencies. For example, if both foo and bar calls another function baz, there is a separate copy of the code and data of baz in each sandbox.
For modules and crates, all publicly accessible functions in the module/crate will be sandboxed in the same way as above, while sharing the same sandbox domain. Note that nested calls to the same module/crate will not enter a new sandbox domain. For example, for a sandboxed crate C with a public function C::foo and a private function C::bar that is called by C::foo, calling C::foo from the trusted domain will enter the sandbox domain of C, while the call to C::bar within C::foo stays in the same sandbox domain as a normal function call. For sandboxed types, all associated methods of the type that are publicly accessible will share the same domain. However, if the transient field is set to true, a fresh sandbox domain will be used for each call to a publicly accessible function of the type/crate.
We assume that different instances in the specification do not overlap with each other in the syntax of the program. For example, if a function foo and a crate c are both specified in the specification, the function foo should not be a member of c. SandCell supports nested sandbox calls so that the code in a sandbox can invoke functions in another sandbox. SandCell could support nested sandbox creation, but we did not find any compelling use case. Thus, our implementation allows only the main sandbox to create new sandboxes.
Specification strategies.
In SandCell, programmers specify sandbox boundaries using syntactic boundaries of program units such as functions and crates. Identifying good isolation boundaries is a well-studied problem: tools such as Privtrans (brumley2004privtrans, ), SeCage (liu2015thwarting, ), PtrSplit (liu2017ptrsplit, ), and Program-mandering (programMandering, ) can automatically compute sandbox boundaries based on information flow analysis or sensitive data annotations. However, these automatic approaches typically enforce fixed kinds of policies (e.g., protecting confidential information). On the other hand, developers with deep knowledge of a codebase understand the program’s behavior and trust relationship and can thus develop policies targeting custom security properties.
Since SandCell is orthogonal to both approaches, it can be combined with automatic partitioning tools or expert-crafted policies. However, even without such tools or expertise, developers can follow simple strategies that provide meaningful security benefits:
  • Library-based (supply-chain): Sandbox entire third-party libraries, especially those with known vulnerabilities or written in unsafe languages. This strategy mitigates supply-chain attacks without requiring understanding of the library’s internals.
  • User-based: In multi-tenant applications, isolate components that process different users’ data from each other. This prevents a compromised user session from attacking other users.
  • Input-oriented: Sandbox components that directly handle external input, such as parsers, decoders, and deserializers. These components are common attack surfaces and benefit from isolation regardless of whether specific vulnerabilities are known.
In our case studies, we found that trust boundaries align naturally with syntactic boundaries (crates, modules, functions), making specification straightforward. We evaluate these strategies in § 8.
5.3. Analysis and Instrumentation
After the user specifies the sandbox boundaries, the next step is to insert necessary code for enforcing sandboxes and passing data between sandboxes. Given the source code, SandCell builds a call graph and a data flow graph of the program. Based on the specification and the call graph, SandCell identifies the public interfaces of types, modules and crates to be sandboxed. Then it inserts calls to the low-level isolation APIs at these interfaces to enforce the isolation at runtime.
For low-level isolation, SandCell adopts an open-sourced isolation library SDRaD by Gülmez et al (gulmez2023dsn, ), and augments it as SDRaD-v2, adding syscall filtering based on zpoline (yasukata23, ) and mimalloc (Leijen2019MimallocFL, ) as the memory allocator. More details can be found in § 7.3. At a high level, SDRaD-v2 provides a set of APIs for managing sandboxes. To switch between sandboxes, SandCell inserts calls to these APIs at the entry and exit points of the sandboxed functions. Also, it provides APIs for management of shared data domains for allocation of objects that two sandboxes can access. More details on the SDRaD-v2 APIs can be found in § 7.
The data passed across sandbox boundaries includes the arguments passed to the sandboxed functions and their return values. To make data from one sandbox available to another sandbox, one approach is to deep copy all such data as in Glamdring (lind2017glamdring, ), PtrSplit (liu2017ptrsplit, ), and GOTEE (ghosn2019secured, ). This approach, however, can lead to a high overhead since deep copying copies all transitively reachable data; it should be avoided when the size of transmitted data is large (gulmezFriendFoeExploring2023, ).
Instead of deep copying, our design copies only stack data and allocate heap data (when it does not contain references to stack data) on a shared data domain that is accessible from both sides of a cross-boundary call. This approach reduces the overhead of copying cross-boundary data. On the other hand, it also requires that cross-boundary data may not be allocated next to each other the relevant memory pages can be assigned to a shared memory region. If we have to copy such data to a shared data domain during cross-boundary calls, we are back to the same problem of high copying cost. To address this, we propose a novel approach to (1) employ static analysis to compute allocation sites that may allocate cross-boundary data in memory, and (2) instrument those allocation sites so that the data is allocated in a shared data domain. This approach avoids the data copying cost by directly allocating cross-boundary data in a shared domain. We detail the static analysis in § 6 and the instrumentation in § 7.
5.4. Run-time Architecture
Figure 3. SandCell run-time architecture.
Figure 3 shows an overview of the SandCell architecture at run time. When the program starts, SDRaD-v2 places the main function and the program stack and heap into a special root domain (➀). When SandCell creates a sandbox (➁), a new SDRaD-v2 domain (➂) is initialized and a new, dedicated, stack and heap area is allocated. Code in the SDRaD-v2 root domain can access data in any new domain created by SandCell, but sandboxed code cannot access the stack or heap of other domains, including the root domain. SDRaD-v2 control data is stored with a separate, monitor domain (➃) that is inaccessible to code in the root or sandboxed domains. The monitor domain will run the code for domain management (➄), the domain-aware memory allocator (➅), and the system call filtering (➆). We will discuss the low-level details in § 7.
Notice that, as shown in Figure 3, crate C is a dependency of both crate A and crate B. However, at run time, because crate A and crate B are sandboxed separately, each of them has its own copy of the code and data of crate C.
5.5. Security Guarantees
SandCell provides memory isolation between sandboxes, preventing memory vulnerabilities from being exploited to access memory outside the sandbox. Suppose that in Figure 3, the attacker exploits a memory vulnerability in crate A and gains the ability to read and write arbitrary memory. Because of the memory isolation provided by SandCell, the attacker can only directly manipulate the memory within Domain #1. Attempts to access memory in other domains, including the root domain, will result in a runtime error. As a practical example, CVE-2018-25032 (CVE-2018-25032, ) is a out-of-bound write triggered by crafted user input in zlib. If sandboxed by SandCell, the out-of-bound write can only modify memory within the sandbox. Another example, CVE-2023-0216 (CVE-2023-0216, ), is an invalid pointer dereferencing vulnerability in openssl. Sandboxing the library confines the invalid memory access within the sandbox.
We note that the guarantee provided by SandCell is typical for sandboxing tools and it does not provide interface safety (SandboxingSok2025SP, ), which requires the data and control flow at the sandbox boundaries be carefully validated to prevent potential misuses or abuse. How to provide interface safety effectively and efficiently in sandboxing is still an area of research.
6. Dataflow Analysis
We next describe the static analysis that identifies which heap objects may cross sandbox boundaries. We first illustrate the problem with a motivating example, then describe our analysis algorithm.
6.1. Problem definition
Goal.
When a sandboxed function is called, any heap data passed as arguments must reside in a memory region accessible to both the sandbox and the main program. Our analysis identifies these cross-boundary heap objects at compile time, so their allocation sites can be instrumented to use the shared memory domain.
Motivating example.
Consider a common pattern where data is accumulated through helper functions, illustrated by the following example, where the main function first loads a series of plugins and then invokes each plugin through a sandboxed call.
fn load_plugins() -> Vec<Plugin> {
let mut plugins = Vec::new(); // Allocation
for path in discover_paths() {
if let Some(p) = parse_plugin(path) {
plugins.push(p); // Container insert
plugins
fn run_sandbox(plugins: &[Plugin]) {
for p in plugins.iter() {
sandbox(p); // Sandboxed call
fn main() {
let plugins = load_plugins();
run_sandbox(&plugins);
The analysis must determine that allocation in load_plugins flows into the sandboxed call in run_sandbox. This requires tracking data through: (1) a return value flowing from load_plugins to main; (2) a function argument flowing from main into run_sandbox; (3) an iterator plugins.iter() that borrows the container; and (4) the push operation that inserts elements into the vector.
Without interprocedural analysis, we cannot connect allocations in load_plugins to uses in run_sandbox. Without modeling container operations, we cannot determine that data pushed into plugins is accessible through iteration. Our analysis handles these cases by building a dataflow graph that is interprocedural and includes summary edges for standard library container methods.
Identifying heap objects.
In Rust, heap allocations are wrapped in safe abstractions like Vec, HashMap, Box, etc. Our analysis recognizes allocations by detecting constructor calls for these standard library containers (e.g., Vec::new(), HashMap::with_capacity()).
6.2. Analysis algorithm
Node Type Description
Local variable A variable declared in a function
Parameter A function’s input parameter
Return value A function’s return value
Field projection A struct field access (e.g., x.field)
Dereference projection A pointer dereference (e.g., *ptr)
Heap allocation A call to Vec::new(), Box::new(), etc.
Edge Type Created For
Assign Direct assignment: x = y
Ref Reference creation: x = &y
Aggregate Struct/tuple construction
CallArg Argument $\to$ parameter at call sites
Return Return value $\to$ destination at call sites
StdSummary Standard library function summaries
Figure 4. Node and edge types in the dataflow graph.
Our analysis operates on Rust’s Mid-level Intermediate Representation (MIR), which the compiler generates before optimization. We construct an interprocedural dataflow graph (DFG) that captures how values flow through the program, then perform backward reachability to find allocation sites.
Dataflow graph structure.
The DFG is a directed graph where:
  • Nodes represent program locations that hold values: local variables, function parameters, return values, struct fields, and heap allocation sites.
  • Edges represent data flow from sources to destinations. An edge from node $A$ to $B$ means the value at $A$ may flow to $B$.
Figure 4 summarizes the node and edge types in our DFG.
We create edges for the following program constructs:
  1. (1)
    Assignments (x = y): Add an Assign edge from y to x.
  2. (2)
    References (x = &y): Add a Ref edge from y to x. We also handle reborrowing patterns like x = &*y, where we first create a projection node for *y, then add a Ref edge to x.
  3. (3)
    Dereferences (x = *y): Add an Assign edge from the projection node *y to x. The dereference node is connected to the pointer’s base via an Aggregate edge.
  4. (4)
    Struct/tuple construction (s = S { f: v }): Add an Aggregate edge from v to the field node s.f, and another Aggregate edge from s.f to s.
  5. (5)
    Function calls (r = f(a, b)): Add CallArg edges from each argument to the corresponding parameter of f, and a Return edge from f’s return value to r.
  6. (6)
    Heap allocations (v = Vec::new()): Create a heap allocation node and add an Assign edge to v.
For standard library functions whose bodies are not available or too complex to analyze, we use pre-defined summaries to specify which arguments flow to the return value. For example, Vec::iter() returns an iterator over its receiver; so we add a StdSummary edge from the receiver to the result.
Handling indirect calls.
Function pointers and trait objects create indirect calls where the target is not statically known. The analysis tracks how function pointers flow through the DFG and uses this information to resolve possible call targets. Critically, DFG construction and call target resolution are performed iteratively: we first build an initial DFG using only direct calls, then resolve indirect call targets based on dataflow, add edges for newly discovered calls, and repeat until no new targets are found. This iterative refinement ensures that the call graph and DFG are mutually consistent.
Backward reachability.
Given the DFG, we find allocation sites by traversing edges backward (from destination to source) starting from sandboxed call arguments. The algorithm is a straightforward worklist-based graph traversal: we initialize the worklist with all arguments of sandboxed function calls, then repeatedly pop a node, follow its incoming edges, and add unvisited source nodes back to the worklist. When a heap allocation node is reached, we record it as a cross-boundary allocation. The traversal continues until the worklist is empty.
Context sensitivity.
The backward traversal is call-site sensitive: when traversing a Return edge from callee to caller, we record the call site on a stack. Later, when traversing a CallArg edge, we check that it matches the recorded call site. This prevents spurious paths that would “return” to a different caller than the one that made the call. The analysis uses a configurable stack depth limit to bound analysis time while maintaining precision for common patterns.
7. Implementation
In this section, we provide the implementation details deferred in the previous sections.
7.1. Program Analysis
SandCell’s analysis is implemented as a plugin to rustc. The compiler translates Rust source code through several levels of intermediate representations (IRs). The Mid-level IR (MIR) is the last IR before translation into the language-neutral LLVM IR. MIR has a much simpler syntax than the surface syntax of Rust but still retains variable type information, which will be stripped away during “lowering”, i.e., translation to LLVM IR. All SandCell analysis are performed at the MIR level.
In MIR, each function has a list of Locals, including the function arguments, temporaries, local variables, and the return value. Each Local might contain a list of Places that correspond to memory locations, e.g., fields of a struct. We map each Place to a node in the dataflow graph and create edges between Places on the left-hand side and right-hand side of an assignment as discussed in § 6. The rest of analysis has been described in § 6.
7.2. Instrumentation
The instrumentation phase inserts code to enforce the sandbox boundaries specified by the user. For each function that crosses the sandbox boundary, a wrapper function is inserted to switch the context to the callee’s sandbox by calling SDRaD-v2 APIs, which will be elaborated in § 7.3. Then the instrumentor redirects the original function call to the wrapper function.
For sharing heap objects, the instrumentor modifies the allocator of the heap object using the standard library’s allocator API. In Rust’s standard library, heap objects are generic over allocators; e.g., Vec<T, A: Allocator> can be instantiated with different allocators. By using Vec<i32, SandCellAlloc>, we direct all memory allocations associated with Vec to our custom allocator SandCellAlloc, which allocates memory in a shared data domain. For types that do not support custom allocators, SandCell wraps their allocations and all the following mutations with a macro that redirects the allocations to the shared data domain.
After the analysis, we statically know how many sandboxes we need in an application and what objects will be shared between which sandboxes. Then for each combination of sandboxes that share data, a shared data domain is allocated and a static ID is assigned to that domain. For example, if the analysis determines that sandbox A, B, and C share data, then a shared data domain is assigned to that combination. [^2] It is worth noting that in our use cases all shared data are between two sandboxes. At allocation sites that allocate objects for a shared data domain, the instrumentor passes the domain’s static ID to SandCell’s SandCellAlloc allocator so that those objects are allocated in that particular data domain.
For stack objects, the instrumentor generates code to copy the stack object to the callee’s stack, following its type information. For references, the instrumentor generates code to copy the referenced object.
7.3. In-Process Isolation
The low-level isolation mechanism used by SandCell is responsible for enforcing the isolation between different sandboxes and providing primitives to transition securely from one sandbox to another. SandCell implements hardware-assisted in-process isolation via the x86-64 processor PKU feature by building on an open-source isolation library, Secure Rewind & Discard (SDRaD(gulmez2023dsn, ), that implements lightweight primitives for in-process isolation and fault recovery. We augment SDRaD with support for a Rust-compatible memory allocator and harden it by incorporating a system call filtering engine. For clarity, we refer in the following description to SDRaD when functionality is provided by the original SDRaD library, and to SDRaD-v2 when describing added functionality.
Secure Rewind and Discard
. Each SDRaD domain must be assigned a unique ID when they are initialized using the sdrad_init API call. SandCell is responsible for assigning an ID to each sandbox, and providing the corresponding ID to the sdrad_enter(id) call that is used to enter an initialized domain. Since SandCell’s isolation boundaries are all function calls, SandCell inserts sdrad_enter(id) calls around sandboxed functions at call sites where SDRaD needs to switch between caller’s and callee’s domains. A corresponding sdrad_exit() call marks the end of code execution, and is used to mark transitions back to the caller’s domain after a sandboxed function returns. SDRaD supports multi-threading by intercepting calls to pthread_create and pthread_exit to manage thread-specific isolation. During a domain switch, SDRaD changes the stack pointer to point to the stack of the domain that is being entered. While executing inside a sandbox, allocations via the POSIX malloc(size) API are redirected to use the heap that belongs to the current SDRaD domain. Allocations that occur outside a domain, but should reside on a specific domain’s heap need to use the sdrad_malloc(id, size) and sdrad_free(ptr) API that take a domain ID as an additional argument. Any memory access to data that is not within a sandbox domain is blocked by SDRaD’s isolation. SDRaD lacks PKU hardening features such as binary rewriting of unauthorized wrpkru instructions or system-call filtering.
Changes in SDRaD-v2
. To ensure that heap-allocated objects are placed into memory belonging to a particular domain, SDRaD provides a domain-aware memory allocator built on the Two-Level Segregated Fit (TLSF) heap allocator (Masmano04, ). However, TLSF implementation used in SDRaD only supports 8-byte alignment for allocations, whereas certain common Rust crates, such as hashbrown, used by the cargo test and cargo bench tools expect allocations to be made at 16-byte alignment. SDRaD-v2 replaces TLSF with the mimalloc general-purpose allocator (Leijen2019MimallocFL, ). In addition, the TLSF-based SDRaD allocator places allocator metadata on the domain heap, where it could be corrupted by memory errors that occur within sandboxed code. The mimalloc allocator, similarly, places allocator metadata in a “backheap” area, that by default would be placed on the domain heap. To harden the SDRaD-v2 allocator, we modified mimalloc to ensure that all metadata is allocated exclusively from the monitor domain, rather than from a domain’s backheap. This guarantees that allocator metadata is restricted to the monitor domain, preventing any modifications from sandboxed code. In SDRaD-v2, the allocator algorithm that is run when malloc or sdrad_malloc is called (➅ in Figure 3), always runs in the monitor domain (➃).
To harden PKU-based isolation, we extend SDRaD with the ERIM binary scanning method to filter out xrstor and wrpkru instructions. In the following section, we explain our approach to hardening using the syscall filtering method.
In total, our changes to SDRaD amount to around $\approx$1.2K lines of C/C++ code, including modifications to mimalloc.
7.4. In-process System Call Filtering.
SandCell needs a solution for system call interposition 1) to harden PKU-based in-process isolation and 2) to reduce the attack surface for each SandCell domain. Early research on hardening PKU-based isolation using system-call interposition can either suffer from high overhead or offer kernel changes or kernel modules (Vahldiek-Oberwagner19, ; Voulimeneas22, ; Schrammel20, ). To address these limitations, we revisited existing system call interposition methods and employed a newer approach, zpoline (yasukata23, ): a technique for the x86-64 architecture to intercept system calls. It replaces the two-byte syscall or sysenter instructions with a call rax instruction of equal length. Since the system call number is stored in the rax register as per the x86-64 call conventions, the call rax causes a control-flow transfer to a memory address corresponding to the system call number, within the first 500 bytes of memory. zpoline populates this range with a series of no-operation instructions, effectively constructing a NOP sled, which eventually leads to a trampoline transferring control to a user-defined code or a syscall interposer’s code. This design ensures that the rewriting of system call instructions cannot fail, providing a reliable method for intercepting and customizing system call behavior in applications. zpoline handles static binaries before main, but it doesn’t support dynamically generated code. To address this, we hook the mprotect call that marks memory as executable and scan the memory region for syscall instructions and rewrite them. This allows zpoline to trace syscalls in dynamically generated code as well. Our system call policy enforcement, implemented on top of zpoline, consists of approximately 300 lines of code C/C++ code in total.
PKU-Hardening.
SDRaD-v2 enforces similar system call filter policies as previous work on PKU-Based isolation primitives (Vahldiek-Oberwagner19, ; Voulimeneas22, ). Using the system call interpositioning primitive, zpoline, (➇ in Figure 3) we trap system calls from any sandbox domains to a system call filtering engine in the monitor domain (➆). The policy engine applies a default filtering policy that rejects any system calls that attempt to circumvent the SDRaD-enforced PKU policy, such as pkey_mprotect calls that attempt to change the PKU keys associated with memory pages, or open family systems calls to /proc/self/mem which would give a sandbox raw access to the memory of the process. SDRaD-v2 also enforces that pages being dynamically mapped as executable do not contain xrstor and wrpkru instructions which could be used to bypass PKU policies or alter the monitor state. To ensure this, the monitor scans such pages for the presence of these instructions, and enforces a Write-XOR-Execute policy on executable pages after inspection. Also, any system call that attempts to modify the zpoline NOP sled area is rejected by our enforcement mechanism. Systems calls from the monitor domain itself are exempt from this policy to enable to SDRaD domain management logic (➄) to manage the sandbox PKU policies properly.
7.5. TCB Size
At runtime, the trusted computing base includes the memory allocator, the SDRaD monitor, and the generated context-switch code. The memory allocator contains $\approx$10k lines of code. The SDRaD monitor contains $\approx$5k lines of code. The generated context-switch code contains around 20 to 100 lines of code for each sandboxed function.
8. Evaluation
We evaluate SandCell along three dimensions: security, usability, and performance. For security, we demonstrate that SandCell mitigates real-world vulnerabilities. For usability, we apply SandCell to real Rust applications and show that isolation policies can be expressed with little specification effort. For performance, the dominant cost is sandbox context switching; so overhead depends on program behavior and boundary placement: frequent crossings amplify overhead, while boundaries on infrequent paths (e.g., one-shot parsing) incur little overhead. To understand how overhead varies across usage patterns, we first measure per-call switching cost in micro-benchmarks against prior tools and the baseline. With that cost in hand, we apply SandCell to diverse real-world applications with different call patterns to quantify end-to-end overhead. Specifically, we answer the following research questions:
  • RQ1 (User Effort): How much effort is required to write specifications for SandCell?
  • RQ2 (Security): Can SandCell effectively mitigate memory safety vulnerabilities?
  • RQ3 (Micro-benchmarks): How does SandCell compare with existing sandboxing tools in controlled benchmarks?
  • RQ4 (Real-world Overhead): What is the runtime overhead when applying SandCell to real-world applications?
Benchmarks.
Across the evaluation, we use two classes of real-world Rust programs: crates with known memory-safety advisories and larger applications/parsing libraries. We randomly selected four Rust crates with memory safety vulnerabilities recorded in the RustSec Advisory Database: transpose (transpose, ), rouille (rouille, ), elf-rs (elfrs, ), and rust-openssl (openssl, ). We also applied SandCell to larger applications and parsing-heavy libraries, including Servo (servo, ), ripgrep (githubGitHubBurntSushiripgrep, ), oq (oq, ), and async-graphql (asyncgraphql, ).
Experiment setup.
We ran all experiments on a 12-core 13th Gen Intel(R) Core(TM) i7-1360P at 2.10 GHz and 64 GiB of RAM. The machine runs Ubuntu 24.04.3 LTS with version 6.12.25 of the Linux kernel. We disable hyperthreading on the CPU to reduce measurement noise.
8.1. RQ1: User Effort for Specification
Application Spec Lines Sandboxed LOC Boundary
transpose 2 671 6
elf-rs 2 1,342 1
rust-openssl 2 49,306 28
rouille 2 10,422 1
Servo 2 134,708 1
ripgrep 2 2,736 8
oq 2 1,767 5
async-graphql 2 2,465 1
Table 1. Specification size, sandboxed LOC, and public API numbers for benchmark applications or libraries. Notice that here we only count the lines of code for the target crate itself, excluding its dependencies. Although at runtime, the sandboxed code may include code from dependencies as well.
We apply the specification strategies described in § 5 to our benchmark applications. Table 1 summarizes the results.
Library-based isolation.
We apply this strategy to three crates with known memory-safety vulnerabilities:
  • transpose is a matrix library with an out-of-bounds write (RUSTSEC-2023-0080 (rustsecRUSTSEC20230080Transpose, )) triggered by crafted input dimensions.
  • elf-rs is an ELF file parser with a heap buffer overflow (RUSTSEC-2022-0079 (RUSTSEC-2022-0079, )) when parsing malformed binaries.
  • rust-openssl wraps OpenSSL, which has had numerous vulnerabilities including CVE-2023-0216 (CVE-2023-0216, ) (invalid pointer dereference) and RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ) (use-after-free).
User-based isolation.
We apply this strategy to two multi-tenant applications:
  • rouille is an HTTP server framework. It uses zlib for compression, which had vulnerabilities including CVE-2018-25032 (CVE-2018-25032, ), CVE-2022-37434 (CVE-2022-37434, ), and CVE-2002-0059 (CVE-2002-0059, ). We isolate each call to the compression function for each user, preventing a compromised request from accessing other users’ data. Identifying this boundary required tracing through 3 functions from the HTTP request handler.
  • Servo is Mozilla’s experimental browser engine. We isolate the JavaScript execution engine for each tab, preventing a malicious script in one tab from accessing data in other tabs. We identified the script execution component and its entry function from the project’s architecture documentation.
While this strategy requires understanding the application’s data flow, the effort remains manageable.
Input-oriented isolation.
We apply this strategy to three applications that process external input:
  • ripgrep is a fast grep tool. We isolate its GlobSet struct, which matches file names against user-provided glob patterns. We identified this by examining the dependency tree, finding the globset crate, and locating its public interface struct.
  • oq is a command-line tool for querying and transforming data in various formats (JSON, YAML, TOML). We isolate the entire oq crate, which parses untrusted input files.
  • async-graphql is a GraphQL server library. We isolate its parsing function, parse_query, which parses untrusted client queries. The identification was straightforward due to the clear function name.
These components directly handle external input and benefit from isolation regardless of whether there are currently known vulnerabilities.
Specification effort.
As shown in Table 1, each application requires only two lines of specification. Once the isolation boundary is identified—whether a crate, struct, or function—specifying it in SandCell is straightforward. The main effort lies in understanding the application’s structure to identify appropriate boundaries, not in writing the specification itself.
8.2. RQ2: Security Evaluation
We evaluate SandCell’s security effectiveness through two complementary approaches: (1) analyzing how SandCell would contain known CVEs in real-world crates, and (2) validating isolation guarantees using proof-of-concept (PoC) exploits derived from cross-language attack patterns.
Analysis of known vulnerabilities.
Our benchmark crates contain known vulnerabilities including RUSTSEC-2023-0080 (rustsecRUSTSEC20230080Transpose, ), CVE-2018-25032 (CVE-2018-25032, ), RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ), and RUSTSEC-2022-0079 (RUSTSEC-2022-0079, ) (out-of-bound writes), as well as RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ), CVE-2002-0059 (CVE-2002-0059, ) and CVE-2023-0216 (CVE-2023-0216, ) (double-free and invalid pointer dereference). By sandboxing the vulnerable function or crate with SandCell, these vulnerabilities would be contained: if triggered, exploits can only access memory inside the sandbox or will be caught by the low-level isolation mechanism of SandCell, which throws a runtime error when memory outside the sandbox is accessed.
PoC-based validation.
To concretely validate SandCell’s isolation guarantees, we developed three proof-of-concept exploits modeling attack patterns from prior work (claPaper, ; mixedBinariesPaper, ), including both pure-Rust and cross-language attacks:
  • Lifetime soundness exploit: A pure-Rust attack exploiting the lifetime expansion soundness hole (Rust issue #25860 (rust25860, )) to forge a mutable reference to an arbitrary address. This models supply-chain attacks where safe Rust can corrupt memory due to compiler unsoundness. When run inside SandCell’s sandbox, cross-domain writes to the monitor cause memory faults handled by SDRaD-v2, leaving the target unchanged.
  • Lifetime violation via C FFI (claPaper, ): Rust allocates a buffer, hands its pointer to C, and C frees it behind Rust’s back; Rust then reads the freed memory, creating a use-after-free. Under SandCell, when the sandbox attempts to free a buffer allocated by the monitor domain, the monitor intercepts and blocks the operation, preventing the attacker from invalidating host memory.
  • Mixed stack corruption (mixedBinariesPaper, ): When Rust and C coexist, Rust stack frames are not protected by SafeStack; a buffer overflow in C can write past a Rust stack buffer and corrupt adjacent control data (e.g., a return address/guard), enabling control-flow hijack. After isolation, the untrusted C runs inside a sandbox domain; so the overflow corrupts only the sandbox’s own stack; the monitor’s control data remains intact, containing the attack’s impact.
Each PoC includes an attack binary demonstrating the exploit without isolation, and an isolated binary showing mitigation under SandCell. In all cases, cross-domain memory accesses from sandbox code into monitor memory result in runtime faults, successfully containing the attack.
8.3. RQ3: Micro-benchmark Comparison
Figure 5. Execution Time of compress() and uncompress() for SandCell, SandCrust, SDRaD-FFI with SDRaD-v1, SDRaD-FFI with SDRaD-v2, Baseline, and Baseline with mimalloc allocator
We measure per-call overhead using the compression library snappy and the image codec libpng, comparing SandCell against SDRaD-FFI (gulmezFriendFoeExploring2023, ) and Sandcrust (lamowskiSandcrustAutomaticSandboxing2017, ).
Benchmark construction.
Each benchmark program executes exactly one sandboxed function call per iteration, measuring total execution time (sandbox overhead plus workload processing). For snappy, we wrap the compress and uncompress functions, invoking each once per iteration while varying the input buffer size (1B to 2MB, in powers of two). For libpng, we wrap decode_png and invoke it once per iteration on different test images. By varying input size, we observe how fixed sandbox overhead amortizes as workload increases. We perform 100 warm-up iterations followed by 50,000 timed iterations per configuration, measuring wall-clock time and computing mean and standard deviation. We reused the public benchmark artifact from SDRaD-FFI (gulmezFriendFoeExploring2023, ) to ensure a fair comparison. Figure 5 presents the overall execution times across different data sizes.
Baseline comparison.
We evaluated both the standard libc allocator and mimalloc as baselines. While the difference is negligible for small buffer sizes, mimalloc provides a clear advantage for larger inputs. Since SandCell uses mimalloc by default, this explains why SandCell outperforms the libc baseline for larger inputs.
SDRaD-FFI comparison.
We measured SDRaD-FFI with its default TLSF allocator and our SDRaD-v2 variant, which enhances compatibility and security. As expected, SDRaD-v2 incurs additional overhead compared to SDRaD-FFI, primarily due to the allocator choice and the syscall interposition mechanism. Although SDRaD-FFI uses in-process isolation, it performs worse than SandCell because its data transfers rely on copying across domains, whereas SandCell reduces this overhead through shared heap access.
Sandcrust comparison.
Sandcrust, which uses processes for isolation, shows worse performance than SandCell due to the overhead of process context switching and inter-process communication.
Summary.
These micro-benchmarks demonstrate that: (1) the heap-sharing optimization of SandCell effectively reduces sandboxing overhead compared to copy-based approaches (SDRaD-FFI); and (2) in-process isolation performs better than process-based isolation (Sandcrust).
8.4. RQ4: Real-world Application Overhead
SandCell-Copy SandCell-Share
Min Max Min Max
rouille 37.1% 89.0% -0.3% 3.0%
elf-rs 12.7% 96.9% 5.4% 15.8%
oq 5.7% 36.0% 8.4% 33.2%
async-graphql 6.8% 15.8% 5.3% 13.3%
ripgrep 5% 13.4% 2.9% 10.3%
rust-openssl 3.9% 17.7% 0.4% 22.2%
Servo - - -1.4% 9.4%
transpose 3.3% 4.4% 3.0% 4.8%
Table 2. SandCell performance overhead compared to the baseline. Missing data indicates the application requires data sharing and cannot be accommodated by the copy mode. Max and min are maximal and minimal overheads under different workloads. For rouille, we use the throughput here. See Appendix B for details.
We evaluate SandCell’s overhead on eight real-world applications and libraries. We compare SandCell-Copy (data copied between domains) and SandCell-Share (heap-sharing optimization). Table 2 summarizes the performance overhead ranges. Detailed performance data can be found in the appendixAppendix B.
Servo.
We run the Dromaeo JavaScript benchmark suite, which executes thousands of JavaScript operations per test category. Each operation triggers a sandbox entry/exit as the JS engine runs in isolation. The high call frequency (hundreds of calls per millisecond) makes this a stress test for sandbox transition cost. Threads in Servo use channels to communicate, which involve heap allocations. We modified the channel implementation by adding 300 lines of code to make it domain-aware, so that SandCell can change their allocation sites to shared domains and allow communication between sandboxes. Copy mode is not applicable because the inter-thread message channel requires real memory sharing. Overall, the max overhead is 9.4%.
rust-openssl.
We use an HTTP library in Rust, attohttpc (attohttpc, ) to implement seven request-level tasks (get, post, session, assets, multipart, redirect, stream). Together they cover basic GET and POST, a session with multiple requests and JSON parsing, asset-style transfers with HEAD and ETag conditional GET, multipart upload of a  256 KB payload, redirect chains, and a streamed JSON response; all responses are consumed and checksummed to keep work on the hot path. These tasks are executed against public endpoints (e.g., www.google.com and httpbin.org) to exercise realistic HTTP client behavior. These program can be found in our artifacts (appendix A). Each TLS handshake and data transfer invokes multiple sandboxed OpenSSL functions. SandCell-Copy introduces 17.7.0% max overhead; SandCell-Share reduces this to 22.2% max.
transpose.
We integrate the sandboxed transpose function into RustFFT, which uses matrix transpositions as part of FFT computation. We vary matrix sizes from $\times64$ to $\times4096$. Each FFT operation triggers one sandbox call, and data transfer for SandCell-Copy scales quadratically with matrix size. For performance measurement, we run the FFT algorithm benchmark that comes with RustFFT. Both SandCell-Copy and SandCell-Share have 3% to 5% overhead because the FFT computation dominates the running time.
rouille.
We run a local rouille server with a configurable thread pool and a single handler that returns a fixed‑size response body, and the response passes through rouille’s content encoding where compression work is done by the sandboxed function, which calls zlib. We then drive that server with wrk. We configure wrk with 8 threads and 200 connections. During the experiement, we run rouille with 1,2 and 4 threads. Each request triggers one sandbox call to the compress function. We measure both latency and throughput. SandCell-Copy introduces 89% max overhead for latency and 77.2% for throughput; SandCell-Share reduces this to 14.9% max and 3% max.
elf-rs.
We build a readelf-like tool that parses ELF binaries and prints its content. We test on different sizes of binaries. Each file triggers one sandbox call to the parser, with data transfer proportional to file size. We measure the running time for each file. SandCell-Copy introduces 96.9% max overhead; SandCell-Share reduces this to 15.8% max.
oq.
The oq benchmark evaluates the oq CLI (a jq-like object query tool) by running a fixed filter over JSON inputs of increasing size. For each sample file (100k–500k JSON) that contains an items array, we execute oq to parse the input and count the array length. The size of data across the boundary for SandCell-Copy is proportional to the input size. For smaller files, the parsing time is dominated by the sandbox transition overhead, therefore, both SandCell-Copy and SandCell-Share show higher overhead (36.0% and 33.2% respectively). When the input size increases, the parsing time dominates, and the overhead decreases (5.7% and 8.4% respectively).
async-graphql.
We setup a GraphQL server using async-graphql with a synthetic dataset (including 2000 entities and relations between them). Then we measure the end-to-end performance with four kinds of queries, covering fragments, pagination/edges, conditional directives, and nested relationships. Each incoming query triggers one sandbox call to parse_query. The parsed AST is returned to the monitor for execution. SandCell-Copy introduces 15.8% max overhead; SandCell-Share reduces this to 13.3% max.
ripgrep.
We use ripgrep to search directory trees with varying numbers of files (64–512). To check pattern matches, the methods of the isolated struct GlobSet are called several times per file , while the array of matched patterns is sent across the boundary. We measure the end-to-end running time. SandCell-Copy introduces 13.4% max overhead; SandCell-Share reduces this to 10.3% max. However, the rest of the overhead is not from the context switching, instead, it is because of the zpoline syscall interception mechanism has non-negligible overhead for the open syscall, which ripgrep invokes frequently when searching files.
Summary.
Overhead correlates with call frequency and data transfer volume. High-frequency call patterns (e.g., rust-openssl) stress sandbox transition costs, while large data transfers (e.g., elf-rs) stress the copying mechanism. SandCell-Share outperforms SandCell-Copy by eliminating data copying overhead in most cases. In some cases, the data size across isolation boundary is small and the noise dominates, so that SandCell-Share performs similar or slightly worse than SandCell-Copy.
9. Related Work
In-process isolation
. Several studies have investigated the compartmentalization of applications based on different underlying primitives, including hardware-assisted in-process isolation (Rivera16, ; Koning17, ; Hedayati19, ; Vahldiek-Oberwagner19, ; Melara19, ; Sung20, ; Lefeuvre21, ; Schrammel20, ; Wang20, ; Voulimeneas22, ; Kirth22, ; Jin22, ; Chen22, ; gulmez2023dsn, ), process-based isolation (lamowskiSandcrustAutomaticSandboxing2017, ), and software-based isolation (Tan17, ). Most of these approaches focus on compartmentalizing applications written in C or C++ (Koning17, ; Hedayati19, ; Vahldiek-Oberwagner19, ; Melara19, ; Sung20, ; Lefeuvre21, ; Schrammel20, ; Wang20, ; Voulimeneas22, ; Jin22, ; Chen22, ; gulmez2023dsn, ). Additionally, several studies specifically address isolating unsafe code from safe code for Rust application (almohriFideliusCharmIsolating2018, ; riveraKeepingSafeRust2021, ; kirthPKRUsafeAutomaticallyLocking2022, ; liuSecuringUnsafeRust2020, ; bangTRustCompilationFramework2023, ). However, those approaches implement conservative compartmentalization policies that label any objects potentially used by untrusted code as unsafe. This can result in edge cases where most allocations are relegated to the unsafe domain, which in turn diminishes the intended benefits of compartmentalization. Also, SDRaD-FFI (gulmezFriendFoeExploring2023, ) and Sandcrust (lamowskiSandcrustAutomaticSandboxing2017, ) focus on providing a finer-grained isolation boundary with manual function annotation, but they do not scale effectively for larger isolation boundaries, such as an entire Rust crate. In contrast, SandCell provides a highly automated solution for sandboxing both safe and unsafe code in Rust applications.
Language-based isolation
. One idea of language-based isolation is to provide language constructs that allow developers to specify isolation boundaries and manage sandboxes. For example, Enclosure (ghosnEnclosureLanguagebasedRestriction2021, ) allows developers to specify capabilities of libraries. GOTEE (ghosn2019secured, ) puts a Go routine into an Intel SGX enclave and automatically extracts sandboxed code and data from the original code. Both systems allow the sandbox boundary based on one existing language construct (libraries or Go Routines). SandCell further pushes this direction by providing flexible boundaries such as crates, functions and types. Also, prior work target Go applications, while SandCell targets Rust applications. Another idea is to provide a sandbox interface in the language or through library, to support sandboxed execution of external code. JavaScript engine V8 (GoogleV8, ) can be used as a virtual machine that runs user code, whose interaction with the outside world is managed by V8. WebAssembly (WebAssembly, ) provides a sandboxed execution environment for web applications by including bounds checking in the semantics of the instruction set. Libsandbox (githubGitHubOpenJudgeSandbox, ) provides APIs in C/C++ and Python for executing and profiling single process programs in a sandbox. However, these tools are designed to execute foreign code in a one-off manner with a high cost of context switching. Furthermore, retrofitting legacy code into these APIs is not desirable for large-code bases. In contrast, SandCell leverages existing syntax boundaries to alleviate the effort needed for compartmentalization.
Automated privilege separation
[…]
10. Discussion & Future Work
In-process isolation.
Like other in-process isolation relying on Intel MPK, SandCell is constrained by the hardware limit of 16 distinct protection domains. As SandCell’s specification and analysis can work independently of the underlying enforcement mechanism, future work could explore replacing MPK with alternative in-process isolation schemes, such as CHERI-based hardware capabilities (cheri, ) or SFI (Tan17, ).
System Call Filtering Alternatives
. As we discuss in Section 7.4, SandCell integrates the zpoline system call interpositioning as part of SDRaD-SandCell to harden PKU-based isolation and to address Challenge IV. While SDRaD-SandCell is fully compatible with conventional system call filtering techniques used in prior PKU-based isolation approaches (Vahldiek-Oberwagner19, ; Voulimeneas22, ), these techniques typically require kernel modifications to improve their efficiency. Without kernel support, they tend to incur significant performance overhead.
System Call Filtering Enhancements
. Currently, SandCell’s isolation only focuses on memory isolation. However, its system call interpositioning mechanism could also be used to enforce more system call policies. For example, we can extend the specification to allow developers to deny certain system call for a particular sandbox. We leave this as future work.
11. Conclusion
We presented SandCell, a novel compiler-based tool that analyzes and instruments Rust code to provide component isolation; we believe that our open-source release of SandCell paves the way for future advancements in securing the Rust ecosystem.
Ethical Considerations
This research adheres fully to established ethical guidelines. It does not result in any tangible harm nor does it infringe upon human rights. All code and content presented in this paper were originally developed by the authors and are free from plagiarism or unauthorized use of external sources.
Generative AI Usage
We used OpenAI Codex to assist in the debugging process of our Rust codebase and to generate our evaluation scripts/workloads. The debugging output can be straightforwardly verified. The evaluation scripts/workloads were reviewed and modified by the authors to ensure correctness. We also use Google’s Gemini to help searching for related work and benchmarks. The search results were further examined by the authors to ensure relevance. For grammar and spelling checking, we used GitHub Copilot.
Appendix A Open Science
The artifacts are available at https://anonymous.4open.science/r/SandCell-Artifact-2D4E. The artifacts include the implementation of the tool, SandCell, and all the benchmarks used in the evaluation, along with scripts to reproduce the results. The instructions for reproducing the results are included in the README file in the artifact repository.
Appendix B Supplementary Measurement
We list all the data from the evaluation below. For readability, we include both the plots and tables for these experiments.
((a)) Running Time of elf-rs
((b)) Running Time of transpose
((c)) Running Time of ripgrep
((d)) Running Time of async-graphql
((e)) Running Time of oq
((f)) Running Time of attohttpc
Figure 6. Performance benchmarks.
Image Baseline SandCrust SandCell Overhead of Overhead of Overhead
Size SandCell SDRaD-FFI SDRaD-FFI
5.5K $160.17$ +183.06% $164.30$ +2.58% $176.28$ +10.06%
63K $2391.45$ +225.99% $2389.73$ -0.07% $2506.35$ +4.80%
378K $5797.16$ +64.21% $5844.68$ +0.82% $5904.48$ +1.85%
874K $8961.68$ +129.44% $8943.76$ -0.20% $9306.31$ +3.85%
Table 3. The Percentage of Overhead for Sandboxing libpng Compared to Baseline.
Matrix Size Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
64 $512.370(22.420\sigma)$ $537.210(21.470\sigma)$ $4.8\%$ $535.010(22.300\sigma)$ $4.4\%$
[…]
96 $1476.160(38.400\sigma)$ $1520.240(36.120\sigma)$ $3.0\%$ $1525.440(54.950\sigma)$ $3.3\%$
112 $1892.640(65.620\sigma)$ $1961.020(45.780\sigma)$ $3.6\%$ $1954.820(64.460\sigma)$ $3.3\%$
128 $2046.130(44.300\sigma)$ $2125.820(57.530\sigma)$ $3.9\%$ $2130.330(45.710\sigma)$ $4.1\%$
Table 4. RustFFT with transpose: Running Time (milliseconds) on Square Matrices of Different Lengths.
Task Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
[…]
post $51.061(5.741\sigma)$ $62.377(21.081\sigma)$ $22.2\%$ $60.089(13.794\sigma)$ $17.7\%$
session $182.633(21.552\sigma)$ $195.319(56.078\sigma)$ $6.9\%$ $200.248(33.824\sigma)$ $9.6\%$
multipart $61.411(4.607\sigma)$ $74.763(12.469\sigma)$ $21.7\%$ $72.036(17.354\sigma)$ $17.3\%$
[…]
stream $54.910(8.434\sigma)$ $61.294(14.557\sigma)$ $11.6\%$ $62.426(17.467\sigma)$ $13.7\%$
Table 5. HTTP Client backed by openssl: Running Time on Different Tasks.
File Count (8MB each) Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
64 $119.476(11.456\sigma)$ $131.738(13.340\sigma)$ $10.3\%$ $135.476(9.656\sigma)$ $13.4\%$
128 $23.320(12.575\sigma) 22$$760(3.558\sigma) 2.92.9\%$ $233.460(1.739\sigma)$ $4.5\%$
[…]
512 $833.026(35.350\sigma)$ $857.149(40.615\sigma)$ $2.9\%$ $874.843(42.189\sigma)$ $5.0\%$
Table 6. ripgrep’s Running Time on Different Number of Files.
File Size Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
47 $9.532(1.680\sigma)$ $0.601(1.442\sigma) 11$$1.2\%$ $10.928(1.347\sigma)$ $14.6\%$
[…]
1192 $22.910(4.491\sigma)$ $24.441(3.216\sigma)$ $6.7\%$ $33.654(2.131\sigma)$ $46.9\%$
5805 $61.925(2.556\sigma)$ $71.721(2.248\sigma)$ $15.8\%$ $121.957(2.723\sigma)$ $96.9\%$
Table 7. Parsing ELF Files with elf-rs Running Time
Pool Size Baseline SandCell-Share SandCell-Copy
Latency (us) Requests/sec Latency (us) Requests/sec Latency (us) Requests/sec
1 8820.0 1584.41 10130.0 1577.41 13830.0 1155.43
2 5030.0 3180.91 4380.0 3191.21 7360.0 2171.07
4 2540.0 6302.12 2610.0 6115.68 4500.0 3335.04
Table 8. rouille Performance across Different Thread Pool Sizes.
Query Baseline SandCell-Copy Overhead of SandCell-Copy SandCell-Share Overhead of SandCell-Share
static.graphql $34.173(3.280\sigma)$ $39.568(5.655\sigma)$ $15.8\%$ $38.572(2.971\sigma)$ $12.9\%$
dashboard.graphql $29.245(0.882\sigma)$ $33.661(3.156\sigma)$ $15.1\%$ $32(0.878\sigma) 13.313.3\%$
batch.graphql $30.781(2.883\sigma)$ $34.585(1.793\sigma)$ $12.4\%$ $34.121(0.750\sigma)$ $10.8\%$
[…]
Table 9. async-graphql Performance Benchmark across Different Query Workloads.
Input Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
100k $13.143(2.519\sigma)$ $17.511(2.717\sigma)$ $33.2\%$ $17.871(3.466\sigma)$ $36.0\%$
200k $22.894(3.975\sigma)$ $24.826(2.458\sigma)$ $8.4\%$ $24.206(1.069\sigma)$ $5.7\%$
336k $33.901(5.077\sigma)$ $37.675(3.065\sigma)$ $11.1\%$ $37.424(3.467\sigma)$ $10.4\%$
400k $37.334(2.568\sigma)$ $41.679(2.327\sigma)$ $11.6\%$ $40.845(3.113\sigma)$ $9.4\%$
500k $45.952(7.905\sigma)$ $50.595(3.499\sigma)$ $10.1\%$ $50.986(3.096\sigma)$ $11.0\%$
Table 10. oq Performance Benchmark for Different Input Size.
Benchmark Overhead of SandCell-Share
Sunspider -0.007%
v8 1.650%
jzlib -1.432%
dom -0.131%
dromaeo 9.360%
Table 11. Servo Average Overhead on each Testsuite of the Dromaeo Benchmark.